OpenStudy (anonymous):

t+2+(7/t)=0 find all solutions and express them in form a+bi

OpenStudy (ybarrap):

t = -1-i sqrt(6) t = -1+i sqrt(6)

OpenStudy (anonymous):

can u explain it to me plz?